OpenStudy (anonymous):

Ready set go :D Solve y = 2x - 4 by the substitution method. 7x – 5y = 14

OpenStudy (bahrom7893):

7x - 10x + 20 = 14 -3x = -6 x=2 y = 0

OpenStudy (lgbasallote):

7x - 5(2x - 4) = 14 7x - 10x + 20 = 14 -3x = -6 x = 2 y = 0 but i lost to bahrom :(

OpenStudy (bahrom7893):

lol i was actually a little distracted in the beginning too lgba haha.. plus it depends on ur internet speed

OpenStudy (anonymous):

Augh why answer Bahroom answer anyways -_-